Grade 11Physical Chemistry

The radii of Na+ and Cl- ions are 95 pm and 181 pm respectively. the edge length of NaCl unit cell is

keerthi
8 Years agoGrade 11

1 Answer

Subhash Kumar Mehta
8 Years ago
NaCl crystal is of fcc type in which Cl- occupy all corner and Na+ occupy all faces so from the structure of fcc we have diagonal of face=181+2(95)+181=552 so sides=[{(552)^2}/2]^(1/2)=390.32pm »Thanks»Regards»Subhash Kumar»St. Xavier`s College, Rnc